Electricity
This term November holidays drove a change in the order of the laboratories. Lab thirteen Ohm's law was moved up one week and optics was pushed back one week. No renumbering was done.
470 Ohms at 1%. Both of the 1% tolerance resistors were beyond 1% away from the color band value, on the order of 2% to 2.5% different.
The 11:00 setup minus the batteries. This worked better and meant less dial twirling. At 8:00 there was too much confusion. Not to mention that there was no schematic diagram put up at 8:00.
And additional battery holders.
The 8:00 session proceeded with groups each using a single multimeter. This was changed to using two multimeters per group and leaving the settings at 20 Volts on one and 200 milliamps on the other.
The morning included an addition to the car collection along the fence line.
470 Ohms at 1%. Both of the 1% tolerance resistors were beyond 1% away from the color band value, on the order of 2% to 2.5% different.
The 11:00 setup minus the batteries. This worked better and meant less dial twirling. At 8:00 there was too much confusion. Not to mention that there was no schematic diagram put up at 8:00.
Jaymar, Franchesca, and Kayla
Moke took over data entry for the middle table. Only three rigs could be supported using two multimeters per rig. At one point a nine volt battery was used which mimics six batteries.
Moki entering data
Comments
Post a Comment